Raptor Corporation is a PHC for 2019 and reports the following items:

Operating profit                                                       $300,000

Which does not include:

Long-term capital gain                                              180,000

Dividends received (25%-owned corporation)         190,000

Interest                                                                       200,000

Gross income                                                                        ?

Salaries expense                                                        (150,000)

General and administrative expense                           (15,000)

Dividends-received deduction                                            (?)

Taxable income                                                                    ?

Regular tax liability                                                              ?

What is the PHC tax for Raptor assuming they pay no dividends to their shareholders?

This is the way the question is asked. Can you please help me understand how to solve?
Answer #1

Operating profit = $300,000

Which does not include:

Long-term capital gain = 180,000

Dividends received (25%-owned corporation) = 190,000

Interest = 200,000

Gross income 300000+180000+190000+200000 = 870,000

Salaries expense (150,000)

General and administrative expense (15,000)

Dividends-received deduction 190000 - (190000*25%) = 142500

Taxable income 870000 - (150000+15000+142500) = 562500

*Regular tax liability = 202,625

Calculation of regular Tax

Tax on first $100,000 of taxable income     =   $ 22,250

Tax on remaining taxable income ($462500×0.39)   = $180,375

Income tax liability    =   $ 202,625

What is the PHC tax for Raptor assuming they pay no dividends to their shareholders?

Answer :

Taxable income = 657,500

Less: federal tax = (202,625)

Less: Net capital gain (net of taxes)* = (109,800)

Plus DRD = 142,500

PHCI = 487,575

PHC Tax (0.20 tax rate×487,575) = 97,515

(* 180,000-(180,000 * 0.39) = 109,800)